# Training towards significance with the decorrelated event classifier transformer neural network

### v1

@@ Abstract
Keywords:
mass resonance search, transformer, significance
1 Introduction
@@
Experimental particle physics is often performed by colliding particles and observing their interaction with detectors. Particles generated from the collisions are investigated by reconstructing the detector data. A common method used in particle searches is to search for a resonance in the reconstructed mass distribution. An example of the resonance is shown in figure 1. Signal events create a peak that can be seen above background events. To estimate the number signal and background events, the mass distribution is fitted with a peaking signal and smooth background function. The fitted mass region is set to be larger than the signal peak width to estimate the background more precisely. This method is referred to as “bump hunting” and was used to search for the Higgs boson CMS-Higgs ; ATLAS-Higgs . The sensitivity of the method can be quantified in terms of significance Significance ; CERN-statistics . A larger significance corresponds to a smaller probability that the bump was created by random statistical fluctuations of the background.
@@
Figure 1: (Left) Mass of reconstructed Higgs boson candidates from the $H\rightarrow Z\left(\ell^{+}\ell^{-}\right)\gamma$ decay, where a bump can be seen due the presence of the Higgs boson particle. The data is generated using MADGRAPH5_aMC@NLO, PYTHIA8, and DELPHES3, where the Higgs boson cross section was scaled up by 100, to make the bump visible. (Right) Mass of reconstructed Higgs boson candidates from the $H\rightarrow Z\left(\ell^{+}\ell^{-}\right)\gamma$ decay with the nominal Higgs boson cross section, where the bump can not be seen due to the background.
@@
To increase the expected significance of the bump hunting method, requirements are applied to create a search region to suppress the background, while preserving the signal. The sensitivity of the analysis can be further enhanced by “binning” the search region in other variables. However, if the binning affects the reconstructed mass distribution of the background, estimating the number of signal events can become difficult. An extreme case would be where the background peaks similarly to the signal. Therefore, a desirable characteristic of binning in the other variables is where it does not affect the shape of background distribution. Binning can be performed using distinguishing features of the signal ATLAS-Higgs or by using machine learning techniques CMS-Higgs ; CMS-HtoMuMu ; CMS-HToZGamma that classify the signal and background events. A common machine learning technique used for this task is the boosted decision tree (BDT).
2 Event classifier transformer neural network
@@
Transformer neural networks Transformer are the leading neural network architecture in the natural language processing field. The architecture has been applied to particle physics in a network called Particle Transformer ParticleTransformer . The network identifies the type of particle that produced a jet, which is a cluster of hadrons whose momentum lie within a cone. In this work, a new transformer-based neural network called event classifier transformer is proposed that classifies signal and background events to bin a bump-hunt analysis.
@@
The architecture of the event classifier transformer is shown in figure 2. The inputs are the features of the event, such as the momenta or angles of particles. To be able to use the transformer architecture, each input or each set of inputs is “embedded” into a token, which is a set of $N_{\text{repr.}}$ numbers, using separate feed-forward neural networks. In contrast to the transformer, there is no positional encoding because event features do not have sequential dependence. The tokens are normalized using layer normalization LayerNormalization and passed to a multi-headed attention layer (MHA) Transformer . The output is summed to employ a residual connection ResidualLayer and normalized with layer normalization. A position-wise feed-forward neural network Transformer is applied with residual connection and layer normalization. The output is a group of tokens, where each token is a contextual representation of the input feature. Following a simplified version of the CaiT ClassToken approach and the Particle Transformer, a trainable class token is passed as a query to an MHA and the contextualized tokens concatenated with the class token is passed as a key and value. The output employs a residual connection and layer normalization and is passed to a linear layer. The linear layer output is a single value that is passed through a sigmoid to represent the probability that the event is from the signal process. To prevent overtraining, dropout layers Dropout are used in the scores of the MHA and before the residual connection during training.
@@
3 Training techniques for enhancing significance
@@
These techniques are described in turn below.
3.1 Specialized loss function with mass decorrelation
@@
A common loss function is the binary cross-entropy (BCE) loss BCE ,
@@
where the minimum corresponds to the condition {linenomath}
@@
In this work, to increase the penalty when $\hat{y}$ and $y$ are different, while keeping the property that the minimum is achieved at $\hat{y}=y$, an alternative loss inspired from BCE called extreme loss is proposed, {linenomath}
@@
For a given value of $\hat{y}-y$, the extreme loss penalizes the neural network more than the BCE loss, as can be seen in figure 3. This extreme loss heavily suppresses the background that have network predictions close to 1. Because the expected significance when binning a bump-hunt analysis is sensitive to backgrounds that have high network predictions, this loss can help increase the significance.
@@
To decorrelate the neural network output with the reconstructed mass, Distance Correlation (DisCo) regularization is used DisCo . DisCo measures the dependence between $\hat{y}$ and mass, where the value of DisCo is zero if and only if $\hat{y}$ and mass are independent. DisCo is multiplied by a factor $\lambda$ and added to the classifier loss function,
$\text{Loss}=\text{Loss}_{\text{classifier}}\left(y,\hat{y}\right)+\lambda\cdot\text{DisCo}\left(\text{mass},\hat{y}\right).$ (5)
@@
3.2 Data scope training
@@
{linenomath}
$\begin{array}[]{ccccc}\text{Loss}&=&\text{Loss}_{\text{\text{classifier}}}\left(\hat{y},y\right)&+&\lambda\cdot\text{DisCo}\left(\text{mass},\hat{y}\right).\\ &&y=\left[0,1\right]&&y=0\\ &&\hat{y}\in\text{narrow mass window}&&\hat{y},\text{mass}\in\text{wide mass window}\end{array}$ (9)
@@
3.3 Significance based model selection
When training networks, the best trained network among the training epochs can be chosen based on an evaluation metric. The loss is often used as the evaluation metric, but the network that has the minimum loss can be different from the network that has the best significance Significance ; LossVsSignificance . In this work, the expected significance is used as a metric to select the best network among the epochs.
@@ 3.3 Significance based model selection
The significance Significance of each bin is calculated with
@@
where $N_{S}$ is the number of signal events and $N_{B}$ is the number of background events within a mass window containing the 5 to 95 percentile of the signal. Note that the mass window width can change depending on the bin. The combination of significance CombinePvalues over the bins is calculated with
@@
4 Example analysis
In this work, a search for the process $H\rightarrow Z\left(\ell^{+}\ell^{-}\right)\gamma$ is considered, where $\ell^{+}\ell^{-}$ represents an $e^{+}e^{-}$ or $\mu^{+}\mu^{-}$ pair. Such a study has been performed by CMS CMS and ATLAS ATLAS with the Run 2 LHC data using a luminosity of around $150~{}\text{fb}^{-1}$. The expected significance for this standard model process of each analysis is $1.2\sigma$ CMS-HToZGamma ; ATLAS-HToZGamma . When adding the expected Run 3 LHC data LHCRun3 , where $250~{}\text{fb}^{-1}$ of data could be collected, and assuming similar analysis sensitivity the expected significance is $2\sigma$. To reach a $3\sigma$ significance, an additional $600\ \text{fb}^{-1}$ of data is required. This will be achieved in the High Luminosity LHC HL-LHC era that is planned to start in 2029, where around $300\ \text{fb}^{-1}/\text{year}$ of data is expected. To reduce the amount of time to obtain evidence $\left(3\sigma\right)$ of this decay, a neural network approach is explored.
@@ 4.1 Data set
The following requirements, which are similar to the CMS analysis CMS-HToZGamma , are applied to the data set:
@@
After applying these requirements, there are 9 million events for the signal and 5 million events for the background.
4.2 Input features for machine learning techniques
@@
  • Lepton identification: Flavor of $\ell$, being either an electron or a muon.
  • Mass reconstruction uncertainty: $\sigma_{m}^{\ell\ell\gamma}$.
  • Kinematic angles: $\eta_{\gamma}$, $\eta_{\ell}^{\text{leading}}$, $\eta_{\ell}^{\text{subleading}}$, minimum $\Delta R\left(\ell^{\pm},\gamma\right)$, maximum $\Delta R\left(\ell^{\pm},\gamma\right)$, production and decay angles $\cos\Theta$, $\cos\theta$, and $\phi$ ProductionDecay1 ; ProductionDecay2 .
  • Momenta: $p_{T}^{\ell\ell\gamma}/m_{\ell\ell\gamma}$ and $p_{Tt}^{\ell\ell\gamma}$ Ptt .
  • Momenta correlated with reconstruction mass: $p_{T}^{\gamma}/m_{\ell\ell\gamma}$, $p_{T}^{\text{leading }\ell}$, and $p_{T}^{\text{subleading }\ell}$. When included in machine learning techniques, such as BDTs, to bin the search region, the background mass distribution in certain bins tend to peak close to the Higgs boson mass, which can be seen in figure 4. This behavior introduces difficulties in estimating the number of signal events, so that these features are typically excluded for the inputs of the machine learning technique. However these features can be used as inputs for neural networks that are trained to be decorrelated with the reconstructed mass.
Figure 4: (Left) Reconstructed $m_{\ell\ell\gamma}$ background distributions, where each histogram is a bin in the XGBoost output distribution with equal number of signal events. Lower signal percentile (sig. p) values correspond to higher output values. $p_{T}^{\gamma}/m_{\ell\ell\gamma}$, $p_{T}^{\text{leading }\ell}$, $p_{T}^{\text{subleading }\ell}$ are not used in the training. (Right) Reconstructed $m_{\ell\ell\gamma}$ background distributions, when training includes $p_{T}^{\gamma}/m_{\ell\ell\gamma}$, $p_{T}^{\text{leading }\ell}$, $p_{T}^{\text{subleading }\ell}$ inputs for XGBoost. For certain bins, the backgrounds peaks close to the Higgs boson mass of 125 GeV, which introduces difficulties in estimating the number of signal events. Correlation represents the magnitude of difference in the shapes between the machine learning bins.
4.3 Machine learning techniques
@@
  • Boosted Decision Trees using the TMVA framework TMVA , which has 850 trees, with a minimum node size of 2.5%.
  • XGBoost XGBoost with 100 number of boosting rounds with a maximum depth of 3.
  • Feed-forward neural network that has an input layer with $N$ nodes, a hidden layer with $4N$ nodes with a $\tanh$ activation function, and an output layer with 1 node with a sigmoid activation function. The network is implemented with PyTorch Pytorch .
  • Deep feed-forward neural network that has an input layer with $N$ nodes and three hidden layers with $3N$, $9N$, and $3N$ nodes with a $\tanh$ activation function. The last hidden layer is connected to a dropout layer with a rate of 0.1 to prevent overtraining. The output layer has a single node with a sigmoid activation function. The network is implemented with PyTorch.
  • Event classifier transformer with $N_{\text{repr.}}=16$. Each event feature is embedded by a separate feed-forward network that has an input layer with one node, a hidden layer with 4 nodes with a GELU activation function, and a output layer with 16 nodes. The MHA has 4 heads and the position-wise feed-forward network has an input layer with 16 nodes, a hidden layer with 64 nodes with a GELU activation function, and a output layer with 16 nodes. The dropout layer has a rate of 0.1. The network is implemented with PyTorch.
@@
When networks are trained with DisCo loss, 3 additional features are used: $p_{T}^{\gamma}/m_{\ell\ell\gamma}$, $p_{T}^{\text{leading }\ell}$, $p_{T}^{\text{subleading }\ell}$. The data scope training method described in section 3.2 is used. The narrow mass window is from 120 to 130 GeV and the wide mass window is from 100 to 180 GeV.
@@
Each neural network is trained for 1200 training epochs with the Adam optimization algorithm Adam with a learning rate of $10^{-3}$ and a batch size of 8192. The best model between the epochs is selected by finding the model with highest significance on the validation data set. To reduce the amount of time in training, the model is evaluated on the validation data set in 5 training epochs intervals for the first 50 epochs and then 10 training epoch intervals for the remaining training epochs. This method allows the best model search to be sensitive to the early epochs, where the evaluation metric can change dynamically.
4.4 Machine learning technique evaluation metrics

### v2

@@ Abstract
I Introduction
@@
Experimental particle physics is often performed by colliding particles and observing their interaction with detectors. Particles generated from the collisions are investigated by reconstructing the detector data. A common method used in particle searches is to search for a resonance in the reconstructed mass distribution. An example of the resonance is shown in Fig. 1. Signal events create a peak that can be seen above background events. To estimate the number signal and background events, the mass distribution is fitted with a peaking signal and smooth background function. The fitted mass region is set to be larger than the signal peak width to estimate the background more precisely. This method is referred to as “bump hunting” and was used to search for the Higgs boson [1, 2]. The sensitivity of the method can be quantified in terms of significance [3, 4]. A larger significance corresponds to a smaller probability that the bump was created by random statistical fluctuations of the background.
@@
Figure 1: (Left) Mass of reconstructed Higgs boson candidates from the $H\rightarrow Z\left(\ell^{+}\ell^{-}\right)\gamma$ decay, where a bump can be seen due the presence of the Higgs boson particle. The Higgs boson cross section was scaled up by 100 to make the bump visible. (Right) Mass of reconstructed Higgs boson candidates from the $H\rightarrow Z\left(\ell^{+}\ell^{-}\right)\gamma$ decay with the nominal Higgs boson cross section, where the bump can not be seen due to the background.
@@
To increase the expected significance of the bump hunting method, requirements are applied to create a search region to suppress the background, while preserving the signal. The sensitivity of the analysis can be further enhanced by “binning”, where the search region is divided into multiple bins according to other variables. However, if the binning affects the reconstructed mass distribution of the background, estimating the number of signal events can become difficult. An extreme case would be where the background peaks similarly to the signal. Therefore, a desirable characteristic of binning in the other variables is where it does not affect the shape of background distribution. Binning can be performed using distinguishing features of the signal [2] or by using machine learning techniques [1, 5, 6] that classify the signal and background events. A common machine learning technique used for this task is the boosted decision tree (BDT).
II Event classifier transformer neural network
@@
Transformer neural networks [7] are the leading neural network architecture in the natural language processing field. The architecture has been applied to particle physics in a network called Particle Transformer [8]. The network identifies the type of particle that produced a jet, which is a cluster of hadrons whose momentum lie within a cone. In this work, a new transformer-based neural network called event classifier transformer is proposed that classifies signal and background events to bin a bump-hunt analysis.
@@
The architecture of the event classifier transformer is shown in Fig. 2. The inputs are the features of the event, such as the momenta or angles of particles. To be able to use the transformer architecture, each input or each set of inputs is “embedded” into a token, which is a set of $N_{\text{repr.}}$ numbers, using separate feed-forward neural networks. In contrast to the transformer, there is no positional encoding because event features do not have sequential dependence. The tokens are normalized using layer normalization [9] and passed to a multi-headed attention layer (MHA) [7]. The output is summed to employ a residual connection [10] and normalized with layer normalization. A position-wise feed-forward neural network [7] is applied with residual connection and layer normalization, where the same position-wise network is applied to each token. The output is a group of tokens, where each token is a contextual representation of the input feature. Following a simplified version of the CaiT [11] approach and the Particle Transformer, a trainable class token is passed as a query to an MHA and the contextualized tokens concatenated with the class token is passed as a key and value. The output employs a residual connection and layer normalization and is passed to a linear layer. The linear layer output is a single value that is passed through a sigmoid to represent the probability that the event is from the signal process. To prevent overtraining, dropout layers [12] are used in the scores of the MHA and before the residual connection during training. Implementation of the network is publicly available at [13].
@@
III Training techniques for enhancing significance
@@
These techniques are described in turn below, where implementation is publicly available at [13].
III.1 Specialized loss function with mass decorrelation
@@
A common loss function is the binary cross-entropy (BCE) loss [14],
@@
where the minimum corresponds to the condition
@@
In this work, to increase the penalty when $\hat{y}$ and $y$ are different, while keeping the property that the minimum is achieved at $\hat{y}=y$, an alternative loss inspired from BCE called extreme loss $E(\hat{y},y)$ is proposed,
@@
For a given value of $\hat{y}-y$, the extreme loss penalizes the neural network more than the BCE loss, as can be seen in Fig. 3. This extreme loss heavily suppresses the background that have network predictions close to 1 and also the signal that have network predictions close to 0. Because the expected significance when binning a bump-hunt analysis is sensitive to backgrounds that have high network predictions, this loss can help increase the significance.
@@
To decorrelate the neural network output with the reconstructed mass, Distance Correlation (DisCo) regularization is used [15]. DisCo measures the dependence between $\hat{y}$ and mass, where the value of DisCo is zero if and only if $\hat{y}$ and mass are independent. DisCo is multiplied by a factor $\lambda$ and added to the classifier loss function,
$\text{Loss}=\text{Loss}_{\text{classifier}}\left(\hat{y},y\right)+\lambda\cdot\text{DisCo}\left(\text{mass},\hat{y}\right).$ (5)
@@
III.2 Data scope training
@@
Loss $=$ $\text{Loss}_{\text{\text{classifier}}}\left(\hat{y},y\right)$
$y=\left[0,1\right]$
$\hat{y}\in\text{narrow mass window}$
$+\lambda\cdot\text{DisCo}\left(\text{mass},\hat{y}\right).$
$y=0$
$\hat{y},\text{mass}\in\text{wide mass window}$
@@
III.3 Significance based model selection
When training networks, the best trained network among the training epochs can be chosen based on an evaluation metric. The loss is often used as the evaluation metric, but the network that has the minimum loss can be different from the network that has the best significance [3, 16]. In this work, the expected significance is used as a metric to select the best network among the epochs.
@@ III.3 Significance based model selection
The significance [3] of each bin is calculated with
@@
where $N_{S}$ is the number of signal events and $N_{B}$ is the number of background events within a mass window containing the 5 to 95 percentile of the signal. Note that the mass window width can change depending on the bin. The combination of significance [17] over the bins is calculated with
@@
IV Example analysis
In this work, a search for the process $H\rightarrow Z\left(\ell^{+}\ell^{-}\right)\gamma$ is considered, where $\ell^{+}\ell^{-}$ represents an $e^{+}e^{-}$ or $\mu^{+}\mu^{-}$ pair. Such a study has been performed by CMS [18] and ATLAS [19] with the Run 2 LHC data using a luminosity of around $150~{}\text{fb}^{-1}$, where the BDT technique was used in binning the search region. The expected significance for this standard model process of each analysis is $1.2\sigma$ [6, 20]. When adding the expected Run 3 LHC data [21], where $250~{}\text{fb}^{-1}$ of data could be collected, and assuming similar analysis sensitivity the expected significance is $2\sigma$. To reach a $3\sigma$ significance, an additional $600\ \text{fb}^{-1}$ of data is required. This will be achieved in the High Luminosity LHC [22] era that is planned to start in 2029, where around $300\ \text{fb}^{-1}/\text{year}$ of data is expected. To reduce the amount of time to obtain evidence $\left(3\sigma\right)$ of this decay, a neural network approach is explored.
@@ IV.1 Data set
The following requirements, which are similar to the CMS analysis [6], are applied to the data set:
@@
After applying these requirements, there are 9 million events for the signal and 5 million events for the background. The large sample helps to reduce overtraining, where the significance metric is sensitive to the sample region that has high classifier scores and low number of background events.
IV.2 Input features for machine learning techniques
@@
  • $\eta_{\gamma}$, $\eta_{\ell}^{\text{leading }\ell}$, $\eta_{\ell}^{\text{subleading }\ell}$: Pseudorapidity angle of the photon and leptons.
  • Minimum $\Delta R\left(\ell^{\pm},\gamma\right)$, Maximum $\Delta R\left(\ell^{\pm},\gamma\right)$: Minimum and maximum $\Delta R$ between leptons and the photon.
  • Flavor of $\ell$: Flavor of lepton used to reconstruct the $Z$ boson, either being an electron or a muon.
  • $p_{T}^{\ell\ell\gamma}/m_{\ell\ell\gamma}$: $p_{T}$ of the reconstructed Higgs boson candidate divided by the mass.
  • $p_{Tt}^{\ell\ell\gamma}$: Projection of the reconstructed Higgs boson $p_{T}$ to the di-lepton thrust axis [29].
  • $\sigma_{m}^{\ell\ell\gamma}$: Mass reconstruction error of the $\ell\ell\gamma$ candidate estimated by binning signal events in $\eta$ and $p_{T}$ for the photon and leptons, and measuring the signal’s mass width for each bin.
  • $\cos\Theta$, $\cos\theta$, $\phi$: Production and decay angles that determine the differential cross section of $H\rightarrow Z\left(\ell^{+}\ell^{-}\right)\gamma$ and $qq\rightarrow Z\left(\ell^{+}\ell^{-}\right)\gamma$ [30, 31].
  • $p_{T}^{\gamma}/m_{\ell\ell\gamma}$, $p_{T}^{\text{leading }\ell}$, $p_{T}^{\text{subleading }\ell}$: Momenta of the photon and leptons.
Many of the features are correlated with the reconstructed Higgs mass. Therefore depending on which features are included in the machine learning technique and which features the machine learning technique prioritizes, the output of machine learning classifier can also be correlated with the mass. Especially when including the momenta of the photon and lepton in certain machine learning techniques, such as BDTs, to bin the search region, the background mass distribution in certain bins tend to peak close to the Higgs boson mass, which can be seen in Fig. 4. This behavior introduces difficulties in estimating the number of signal events, so that these features are typically excluded for the inputs of the machine learning technique. However these features can be used as inputs for neural networks that are trained to be decorrelated with the reconstructed mass.
Figure 4: (Above) Reconstructed $m_{\ell\ell\gamma}$ background distributions, where each histogram is a bin in the XGBoost output distribution with equal number of signal events. Lower signal percentile (sig. p) values correspond to higher output values. $p_{T}^{\gamma}/m_{\ell\ell\gamma}$, $p_{T}^{\text{leading }\ell}$, $p_{T}^{\text{subleading }\ell}$ are not used in the training. (Below) Reconstructed $m_{\ell\ell\gamma}$ background distributions, when training includes $p_{T}^{\gamma}/m_{\ell\ell\gamma}$, $p_{T}^{\text{leading }\ell}$, $p_{T}^{\text{subleading }\ell}$ inputs for XGBoost. For certain bins, the backgrounds peaks close to the Higgs boson mass of 125 GeV, which introduces difficulties in estimating the number of signal events. Correlation represents the magnitude of difference in the shapes between the machine learning bins.
IV.3 Machine learning techniques
@@
  • Boosted Decision Trees using the TMVA framework [32], which has 850 trees, with a minimum node size of 2.5%.
  • XGBoost [33] with 100 number of boosting rounds with a maximum depth of 3. The BDT was optimized by modifying the number of boosting rounds, where BDTs with more than 100 round have higher overtraining, while other performance metrics are similar.
  • Feed-forward neural network that has an input layer with $N$ nodes corresponding to the number of input features, a hidden layer with $4N$ nodes with a $\tanh$ activation function, and an output layer with 1 node with a sigmoid activation function. The network is implemented with PyTorch [34].
  • Deep feed-forward neural network that has an input layer with $N$ nodes corresponding to the number of input features and three hidden layers with $3N$, $9N$, and $3N$ nodes with a $\tanh$ activation function. The last hidden layer is connected to a dropout layer with a rate of 0.1. The output layer has a single node with a sigmoid activation function. The network is implemented with PyTorch. The network was optimized by modifying the dropout rate, where the network with no dropout had significant overtraining and poorer performance, while the network with a dropout rate of 0.2 had similar performance with the dropout 0.1 network.
  • Event classifier transformer with $N_{\text{repr.}}=16$. Each event feature is embedded by a separate feed-forward network that has an input layer with one node, a hidden layer with 4 nodes with a GELU activation function, and a output layer with 16 nodes. The MHA has 4 heads and the position-wise feed-forward network has an input layer with 16 nodes, a hidden layer with 64 nodes with a GELU activation function, and a output layer with 16 nodes. The dropout layer has a rate of 0.1. The network is implemented with PyTorch. The network was optimized by modifying the dropout rate, where the network with no dropout showed overtraining and poorer performance, while the network with a dropout rate of 0.2 had slighter worse performance than the dropout 0.1 network.
@@
When networks are trained with DisCo loss, 3 additional features are used: $p_{T}^{\gamma}/m_{\ell\ell\gamma}$, $p_{T}^{\text{leading }\ell}$, $p_{T}^{\text{subleading }\ell}$. The data scope training method described in section III.2 is used. The narrow mass window is from 120 to 130 GeV and the wide mass window is from 100 to 180 GeV.
@@
Each neural network is trained for 1200 training epochs with the Adam optimization algorithm [35] with a learning rate of $10^{-3}$ and a batch size of 8192. The best model between the epochs is selected by finding the model with highest significance on the validation data set. To reduce the amount of time in training, the model is evaluated on the validation data set in 5 training epochs intervals for the first 50 epochs and then 10 training epoch intervals for the remaining training epochs. This method allows the best model search to be sensitive to the early epochs, where the evaluation metric can change dynamically.
IV.4 Machine learning technique evaluation metrics